OpenStudy (anonymous):

1. Let's say that we have a rectangle with sides called a and b. If a = 11 inches and b = 3 inches, what is the perimeter of this rectangle in inches

OpenStudy (anonymous):

P=2a+2b =2*11+2*3 =28 inches